40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Dec 11 15:01:57 2019, Jon, Summary, PSL, PMC cavity ringdown measurement IMG_0101.jpg
    Reply  Thu Dec 12 19:20:43 2019, Yehonathan, Update, PSL, PMC cavity ringdown measurement IMG_0105.jpegTEK00001.PNG20191212_151642.jpg
       Reply  Fri Dec 13 12:28:43 2019, Yehonathan, Update, PSL, PMC cavity ringdown measurement Ringdown_InitialProcess.pdf
          Reply  Tue Dec 17 20:45:30 2019, rana, Update, PSL, PMC cavity ringdown measurement 
          Reply  Fri Dec 27 15:01:02 2019, Yehonathan, Update, PSL, PMC cavity ringdown measurement PDAOMResponse.pdfPMCTransmission.pdfRingdownsAndRefs.pdfTwoExponentialFitAOM0.3V.pdf
             Reply  Tue Dec 31 03:03:02 2019, gautam, Update, PSL, PMC cavity ringdown measurement 
       Reply  Mon Dec 16 18:19:42 2019, shruti, Update, PSL, PMC cavity ringdown measurement : beat-note disruption IMG_0040.jpg
Message ID: 15096     Entry time: Thu Dec 12 19:20:43 2019     In reply to: 15093     Reply to this: 15097   15098
Author: Yehonathan 
Type: Update 
Category: PSL 
Subject: PMC cavity ringdown measurement 

{Yehonathan, Rana, Jon}

To check whether we laser is being shut fast enough for the ringdown measurement we put a PD55 in the path that leads to the beat note setup. The beam is picked off from the back steering mirror after AOM and before the PMC.

@Shruti the PD is now blocking the beam to your setup.

As before, we drive the AOM to deflect the beam. The deflected beam is coupled to the PMC cavity. We lock the PMC and then shut the beam by turning off the output of the function generator that provides voltage to the AOM driver.

We measure the transmitted light of the PMC together with the light that is picked off before the PMC. In Attachment 1, the purple trace is the PMC transmission, the green trace is the peaked-off beam and the yellow trace is the function generator signal.

Rana was pointing out that the PDs, the function generator and the scope were not carefully impedance matched, which could lead to erroneous measurements. He also mentioned that the backscattered beam was too bright which might indicate that the PMC is oscillating. To remedy this we lowered the gain of the PMC lock to ~8.

We repeat the measurement after setting all the components to 50ohm (attachment 2). We then realize that the BNC T junction connected on the function generator is splitting the signal between the 50ohm AOM driver and 1Mohm oscilloscope channel which causes distortions as can be seen. We remove the T junction and get a much cleaner measurement (see next elog).

 

It seems like either the shutting speed or the PDs are only slightly faster than the PMC. I also check the AOM driver RF output fall time doing the same kind of measurement (attachment 3).

We suspect the PDs' bandwidth is to blame (although they are quoted to have 10MHz bandwidth).

In any case, this is fast enough for the IMC and arm cavities whose lifetime should be much longer than the PMC.

I will post an elog with some numbers tomorrow.

Attachment 1: IMG_0105.jpeg  1.999 MB  | Hide | Hide all
IMG_0105.jpeg
Attachment 2: TEK00001.PNG  14 kB  | Hide | Hide all
TEK00001.PNG
Attachment 3: 20191212_151642.jpg  4.397 MB  | Hide | Hide all
20191212_151642.jpg
ELOG V3.1.3-